OpenStudy (anonymous):

find the sum to the required number of terms for the geometric series 2-1+1/2 -1/4+... to 6 terms just want to check if i got the answer right. using the formula s=b1*(1-q^n)/(1-q) , i got 21/16 is that right?

Parth (parthkohli):

\[1 + 1/2 - 1/4 + 1/8 - 1/16 \\ = 1+1/4 + 1/8 - 1/16\\ = 1 + 3/8 - 1/16\\ = 1 + 5/8\]

Parth (parthkohli):

So you're correct!

Parth (parthkohli):

When I don't have a paper, I just add those up. Haha

OpenStudy (anonymous):

but 1+5/8 is 13/8

OpenStudy (anonymous):

13/8 = 1.625 and 21/16=1.31

OpenStudy (goformit100):

Where is the lamp @AravindG

Parth (parthkohli):

grr, I meant \(1 + \dfrac{5}{16}\)

OpenStudy (anonymous):

ohh lol ok that makes sense

OpenStudy (goformit100):

grr nice peculiure sound you make

OpenStudy (goformit100):

@ParthKohli it was for you baby